OpenStudy (anonymous):

Using each of the digits 1,2,3,4,5 and 6 only once, find two 3-digit numbers whose product is as large as possible. Explain your answer

OpenStudy (anonymous):

Basically, we are asked to pick two 3 digit numbers, which when multiplied together are a max for this set. . if we try to maximize the hundreds digit: 100*200=20,000 300*400=120,000 500*600=300,000,,this suggests that the hundreds digits are 5 &6 . working on the 10's digit: 610*520=317,200 620*510=316,200 630*540=340,200,,,,,,max,,,lets use these 640*530=339,200 . working on the tens digits: 631*542=342,002,,,,,,max,,,,,answer 632*541=341,912 . Therefore using only the digits 1 to 6, and each only once, the max product is 342,002, from 631*542

OpenStudy (anonymous):

thank you. :3